Parth (parthkohli):

Note: This is NOT a question, but a mind-blowing tutorial. What were slopes and intercepts, btw? Well, right place is what I should say.

Parth (parthkohli):

Well, what are equations in slope-intercept form? They are basically the equations that have a straight-line graph and are in the form: \(\Huge \color{MidnightBlue}{\Rightarrow y = mx + b }\) I hope it was readable :P In the equation, as you see, m is the slope, and b is the y-intercept. If you are given the equation \(\rightarrow y = 3x + 6\), then compare it with y = mx + b. The slope is m i.e., 3 in this case. The y - intercept is b, which is +6 in this case. \(\Huge \color{MidnightBlue}{\Rightarrow SLOPE }\) What is slope? Slope equals change in y over change in x. Another way to think about it is: \(\Large \color{MidnightBlue}{\Rightarrow Slope = {y_{2} - y_{1} \over x_{2} - x_{1}} }\) If you are given two pairs of co-ordinates, then see the y-value in the second pair and subtract that by the y-value in the first pair. Divide that by the difference of \(x_{2} \text{ and } x_{1} \). What is the slope of a line that passes through points (3,4) and (4,5) Well, try to solve that as you know now how to get it. \(\Huge \color{MidnightBlue}{\Rightarrow Intercepts }\) What is an intercept? Intercept is a value of a variable obtained when you set up the other variable as 0. To obtain the y-intercept, set x = 0. To obtain the x-intercept, set y = 0. For example, the equation is given as follows and you have to determine the x-intercept and y-intercept. \(\Large \color{Orange}{\Rightarrow y = {1 \over 3}x + 6 }\) y-intercept would be 6, because when we set x = 0, then it simply cancels out and we are left with y = 6. x-intercept would be -18, because when we set y = 0 and subtract 6 from both sides, we are left with: \(\Large \color{MidnightBlue}{\Rightarrow -6 = {1x \over 3} }\) \(\Large \color{MidnightBlue}{\Rightarrow x = -6(3) = -18 }\) Next tutorial in the series: How to graph when given the slope.

OpenStudy (angela210793):

i still can't read anything [Math Processing Error] :P

Parth (parthkohli):

It's just a bug. Refresh the page.

OpenStudy (anonymous):

Nice and informative...and simple:D

OpenStudy (lgbasallote):

lol jazy...fast reader :D

OpenStudy (angela210793):

lol @Lgb :D

OpenStudy (anonymous):

I pride myself in that lol:P love to read and math is somewhat entertaining...nice mix:)

OpenStudy (lgbasallote):

nice @ParthKohli getting better :D one bad comment though...you still cant estimate the size of the latex...some text needs to be smaller...some need to be bigger

Parth (parthkohli):

Lol it was fully intentional.

OpenStudy (lgbasallote):

\[\large \frac{y_2 - y_1}{x_2 - x_1}\] is barely readable :P

OpenStudy (anonymous):

share... is their are people who is numb?

OpenStudy (anonymous):

The way y always thought of y=mx+b is if you are given a linear equation not in y=mx+b is to just solve for y.

OpenStudy (anonymous):

*I